# Relative_Constructibility_via_Restricted_ASMs.bib

### v1

@@ 3.3 Associations with Constructibility
We use $L[A]$ to denote the constructible universe relative to a set $A$, and $L$ to denote $L[\emptyset]$. For each ordinal $\alpha$, let $L_{\alpha}[A]$ respectively denote the $\alpha$-th level $L[A]$, so $L_{\alpha}[\emptyset]$ is just $L_{\alpha}$. The reader may refer to Chapter 13 of [jech] for more details.
@@
Let $\kappa$ be a limit ordinal. We say $B$ is $\kappa$-computable with parameters from $A$ with short runs, or $B\leq^{P,s}_{\kappa}A$, if there is an RASMP $\mathfrak{M}$ witnessing $B\leq^{P}_{\kappa}A$ and a short terminating run of $\mathfrak{M}$ with input $A$ and output $B$. Here, $\mathfrak{M}$ is said to witness $B\leq^{P,s}_{\kappa}A$.
@@
For a limit ordinal $\alpha$, an $\alpha$-machine according to Koepke (see [koepke2]) is essentially a Turing machine with tape length $\alpha$, oracle length up to $\alpha$, and finitely many fixed ordinal parameters $<\alpha$. We can think of an input to an $\alpha$-machine as a code $\langle X,O\rangle$ of a pair of sets, where $X$ represents the non-oracle component of the input and $O$ represents the oracle. At limit steps the limit inferiors of previous head and state positions, along with the previous contents of each cell, are taken. The machine halts iff it terminates in $<\alpha$ many steps.
@@
Let $\alpha$ be a limit ordinal. An $\alpha$-machine $M$ can be simulated by an RASMP $\mathfrak{M}$ in the sense that for any $A,B\subset\alpha$, $M$ halts on input $A$ with output $B$ iff $\mathfrak{M}$ has a short terminating run with input $A$ and output $B$.
@@
It is not difficult to check that modifying Example 2.29 by substituting $\alpha$ for $\omega$, and adding oracle reads based on how the oracle is coded into the input (through e.g. adding a third tape), works out just fine. The finitely many ordinal parameters of an $\alpha$-machine can be ported wholesale to an RASMP as the latter’s ordinal parameters; the way an RASMP determines the limit states of a computation also perfectly mimicks how limit configurations of an $\alpha$-machine are defined. Moreover, an $\alpha$-machine halting on input $A$ with output $B$ is equivalent to the RASMP simulating it having a short terminating run with input $A$ and output $B$. ∎
Definition 3.33 (Koepke, [koepke2]).
@@
Let $\mathfrak{M}$ be an RASMP with parameter set $K$ witnessing $B\leq^{P,s}_{\kappa}A$, and $Sq$ be a short terminating run of $\mathfrak{M}$ with input $A$ and output $B$. Suppose we want to find out whether some $\beta<\kappa$ is a member of $B$. To do so, we draw up a dependency tree of terms of which interpretations among members of $ran(Sq)$ are needed to decide if $\beta\in B$. This tree is finitely branching by the bounded exploration postulate, and it has number of levels no more than $dom(Sq)$. By the regularity of $\kappa$, it has size $<\kappa$.
Now, the statement $x\in B$ is semantically equivalent over $V$ to a $\Sigma_{1}$ formula with free variable $x$ and parameters among $K\cup\{dom(Sq)\}$ saying that such a (unique) tree of terms and interpretations exist. The construction of said tree, due to its simplicity, can be carried out correctly in $L[A]$ through transfinite recursion. Moreover, $L[A]$ recognises the tree has size $<\kappa$, so indeed $B\in\mathbf{\Sigma_{1}}(L_{\kappa}[A])$.
We can define an RASMP $\mathfrak{M}^{\prime}$ that copies what $\mathfrak{M}$ does on each input until $\mathfrak{M}$ terminates, and flips every bit of the output tape thereafter. Now $\mathfrak{M}^{\prime}$ witnesses $\kappa\setminus B\leq^{P,s}_{\kappa}A$, so by the argument in the preceding paragraph we have $\kappa\setminus B\in\mathbf{\Sigma_{1}}(L_{\kappa}[A])$. ∎
Definition 3.38 (Koepke, [koepke2]).
@@
By Fact 3.39, $B$ is $\kappa$-computably enumerable in $A$, as witnessed by some $\kappa$-machine $M$. Let $\mathfrak{M}$ be an RASMP simulating $M$ as per Proposition 3.32. We define an RASMP $\mathfrak{M}^{\prime}$ that, when given input $A$, dovetails and runs $M$ through $\mathfrak{M}$ on each input in
@@
for arbitrarily many steps $<\kappa$, in order to determine if $M$ halts on said input. In the end, $\mathfrak{M}^{\prime}$ should consolidate the findings and output them as a subset of $\kappa$.
More formally, add to $\sigma(\mathfrak{M})$ fresh constant symbols $\ulcorner c_{0}\urcorner$, $\ulcorner c_{1}\urcorner$, $\ulcorner c_{2}\urcorner$, $\ulcorner d\urcorner$ and a fresh unary function symbol $\ulcorner R\urcorner$, and have the result be $\sigma(\mathfrak{M}^{\prime})$. Set all constants to 00 and $R$ to be empty by default. We are left to describe the transition function of $\mathfrak{M}^{\prime}$, which we shall do by cases. In our description, $I$ denotes the input to $\mathfrak{M}^{\prime}$.
@@ 2. Case 2:
    $c_{0}\neq 0\wedge d\neq 0$”. We do the dovetailing. In other words, for each $\beta$ satisfying “$\neg R(\beta)\wedge\beta\leq c_{0}$”, we use $\mathfrak{M}$ to simulate $M$ on input $\langle\{\beta\},I\rangle$ for up to $c_{0}$ steps, using $c_{1}$ and $c_{2}$ to keep track of the counts. At the end of every step in simulation, we check if $\mathfrak{M}$ terminates. If so, we set $R(\beta)$ to true and move on to simulate $M$ on input $\langle\{\beta+1\},I\rangle$, if $\beta\neq c_{0}$. At the end of all the simulations and possible updates to $R$, set $c_{1}$ and $c_{2}$ to 00 and increment $c_{0}$. The constant $d$ is left unchanged.
@@ 4. Case 4:
  • $c_{0}$ first returns to 00 in a state other than an initial state after exactly $\kappa+1$ many steps, at which point $M$ would have been simulated on every input in
@@
    for up to arbitrarily many steps $<\kappa$, with all eventual haltings recorded in $R$.
@@
Lemmas 3.37 and 3.40 reflect the difference in computability between considering all terminating runs of RASMPs and considering only short terminating runs, for regular ordinal tape lengths. In tandem, the lemmas also give us a nice characterisation of the reducibility relation $\leq^{P}$.
Lemma 3.41.
@@
Apply Lemma 3.41 with $\emptyset$ in place of $A$. ∎
@@ 4 Discussion and Questions
  • a “high-level” machine model of generalised computation such as RASMPs, and
  • an algebraic model of generalised computation such as relativised constructibility
@@
Just as practical implementations of finitary ASMs do not aim to model algorithmic intuition beyond the limits of Church-Turing thesis, RASMPs should not allow for algorithms that are non-constructible. We can view RASMPs as a paradigm for easy high-level descriptions of algorithms on arbitrary sets, one that gives the user enough structure to organise their ideas, and serves as guardrails to rule out clearly non-algorithm descriptions.
@@
    $\mathsf{ZFC}+\text{``}\exists\text{ a transitive model of }\mathsf{ZFC}\text{''}\vdash\text{``}&\exists\text{ a transitive model of }\mathsf{ZFC}\ +\\ &\text{`}V\neq L[A]\text{ for all }A\in V\text{'''.}\end{aligned}$
@@
    $\mathsf{ZFC}+\text{``}\exists\text{ a strongly compact cardinal''}\vdash\text{``}V\neq L[A]\text{ for all }A\in V\text{''}$ ”.
In particular, (2) of Fact 4.3 is implied by Exercise 20.2 of [jech]. Note that $V\neq L[A]$ for all $A\in V$ is equivalent to $(\mathcal{D}_{P},\leq_{\mathcal{D}_{P}})$ having no maximal element, so Fact 4.3 actually gives us
@@
    $\mathsf{ZFC}+\text{``}\exists\text{ a transitive model of }\mathsf{ZFC}\text{''}\vdash\text{``}&\exists\text{ a transitive model of }\mathsf{ZFC}\ +\\ &\text{`}(\mathcal{D}_{P},\leq_{\mathcal{D}_{P}})\text{ has no maximal element''',}\end{aligned}$
@@
    $\mathsf{ZFC}+\text{``}\exists\text{ a strongly compact cardinal''}\vdash\text{``}(\mathcal{D}_{P},\leq_{\mathcal{D}_{P}})\text{ has no maximal element''}$”.
@@
Let us now turn our attention to the bounded relations $\leq^{P,s}_{\alpha}$ and $\leq^{P}_{\alpha}$. Let $\leq_{\alpha}$, $\leq_{w\alpha}$ and $\preceq_{\alpha}$ respectively denote the relations “$\alpha$-recursive in”, “weakly $\alpha$-recursive in” and “$\alpha$-computable in”. The first two relations are well-studied in the once-bustling field of $\alpha$-recursion theory (see e.g. Part C of [sacks]), whereas the last relation comes from the emerging field of $\alpha$-computability theory. It is known (Section 3 of [koepke2]) that
@@
for admissible ordinals $\alpha$. The last inclusion cannot be reversed in general because $\leq^{P}_{\alpha}$ is transitive for all admissible $\alpha$ but $\preceq_{\alpha}$ is not. On the other hand, it is clear that $\leq^{P,s}_{\alpha}\ \subset\ \leq^{P}_{\alpha}$ for all limit $\alpha$, and the reverse is not true due to Lemmas 3.37 and 3.40. Further, for admissible $\alpha$, $\leq_{w\alpha}\ \neq\ \leq^{P,s}_{\alpha}\ \neq\ \preceq_{\alpha}$ in general because $\leq^{P,s}_{\alpha}$ is transitive unlike the other two relations. We know $\leq^{P,s}_{\alpha}\ \neq\ \leq_{\alpha}$ in general because $\leq^{P,s}_{\alpha}$ is upward absolute in the sense of Proposition 3.12, but $\leq_{\alpha}$ is not by Theodore Slaman’s answer in [slaman].
Table 3: Comparing two of our relative computability relations with more standard ones in the of study generalised computability, at arbitrary admissible ordinals $\alpha$.
@@
Appears $\dots$ $\alpha$-recursion $\alpha$-computability Def. 3.8 Def. 3.26

### v2

@@ 3.3 Associations with Constructibility
Let us begin this subsection with a recapitulation of selected set-theoretic concepts.
@@
We use $L[A]$ to denote the constructible universe relative to a set $A$, and $L$ to denote $L[\emptyset]$. For each ordinal $\alpha$, let $L_{\alpha}[A]$ respectively denote the $\alpha$-th level $L[A]$, so $L_{\alpha}[\emptyset]$ is just $L_{\alpha}$. The reader may refer to Chapter 13 of [12] for more details.
@@
Let $\kappa$ be a limit ordinal. We say $B$ is $\kappa$-computable with parameters from $A$ with short runs, or $B\leq^{P,s}_{\kappa}A$, if there is a GSeqAP $\mathfrak{M}$ witnessing $B\leq^{P}_{\kappa}A$ and a short terminating run of $\mathfrak{M}$ with input $A$ and output $B$. Here, $\mathfrak{M}$ is said to witness $B\leq^{P,s}_{\kappa}A$.
@@
For a limit ordinal $\alpha$, an $\alpha$-machine according to Koepke (see [14]) is essentially a Turing machine with tape length $\alpha$, oracle length up to $\alpha$, and finitely many fixed ordinal parameters $\alpha$. We can think of an input to an $\alpha$-machine as a code $\langle X,O\rangle$ of a pair of sets, where $X$ represents the non-oracle component of the input and $O$ represents the oracle. At limit steps the limit inferiors of previous head and state positions, along with the previous contents of each cell, are taken. The machine halts iff it terminates in $\alpha$ many steps.
@@
Let $\alpha$ be a limit ordinal. An $\alpha$-machine $M$ can be simulated by a GSeqAP $\mathfrak{M}$ in the sense that for any $A,B\subset\alpha$, $M$ halts on input $A$ with output $B$ iff $\mathfrak{M}$ has a short terminating run with input $A$ and output $B$.
@@
It is not difficult to check that modifying Example 2.29 by substituting $\alpha$ for $\omega$, and adding oracle reads based on how the oracle is coded into the input (through e.g. adding a third tape), works out just fine. The finitely many ordinal parameters of an $\alpha$-machine can be ported wholesale to a GSeqAP as the latter’s ordinal parameters; the way a GSeqAP determines the limit states of a computation also perfectly mimicks how limit configurations of an $\alpha$-machine are defined. Moreover, an $\alpha$-machine halting on input $A$ with output $B$ is equivalent to the GSeqAP simulating it having a short terminating run with input $A$ and output $B$. ∎
Definition 3.33 (Koepke, [14]).
@@
Let $\mathfrak{M}$ be a GSeqAP with parameter set $K$ witnessing $B\leq^{P,s}_{\kappa}A$, and $Sq$ be a short terminating run of $\mathfrak{M}$ with input $A$ and output $B$. Suppose we want to find out whether some $\beta<\kappa$ is a member of $B$. To do so, we draw up a dependency tree of terms of which interpretations among members of $ran(Sq)$ are needed to decide if $\beta\in B$. This tree is finitely branching by the bounded exploration postulate, and it has number of levels no more than $dom(Sq)$. By the regularity of $\kappa$, it has size $\kappa$.
Now, the statement $x\in B$ is semantically equivalent over $V$ to a $\Sigma_{1}$ formula with free variable $x$ and parameters among $K\cup\{dom(Sq)\}$ saying that such a (unique) tree of terms and interpretations exist. The construction of said tree, due to its simplicity, can be carried out correctly in $L[A]$ through transfinite recursion. Moreover, $L[A]$ recognises the tree has size $\kappa$, so indeed $B\in\mathbf{\Sigma_{1}}(L_{\kappa}[A])$.
We can define a GSeqAP $\mathfrak{M}^{\prime}$ that copies what $\mathfrak{M}$ does on each input until $\mathfrak{M}$ terminates, and flips every bit of the output tape thereafter. Now $\mathfrak{M}^{\prime}$ witnesses $\kappa\setminus B\leq^{P,s}_{\kappa}A$, so by the argument in the preceding paragraph we have $\kappa\setminus B\in\mathbf{\Sigma_{1}}(L_{\kappa}[A])$. ∎
Definition 3.38 (Koepke, [14]).
@@
By Fact 3.39, $B$ is $\kappa$-computably enumerable in $A$, as witnessed by some $\kappa$-machine $M$. Let $\mathfrak{M}$ be a GSeqAP simulating $M$ as per Proposition 3.32. We define a GSeqAP $\mathfrak{M}^{\prime}$ that, when given input $A$, dovetails and runs $M$ through $\mathfrak{M}$ on each input in
@@
for arbitrarily many steps $\kappa$, in order to determine if $M$ halts on said input. In the end, $\mathfrak{M}^{\prime}$ should consolidate the findings and output them as a subset of $\kappa$.
More formally, add to $\sigma(\mathfrak{M})$ fresh constant symbols $\ulcorner c_{0}\urcorner$, $\ulcorner c_{1}\urcorner$, $\ulcorner c_{2}\urcorner$, $\ulcorner d\urcorner$ and a fresh unary function symbol $\ulcorner R\urcorner$, and have the result be $\sigma(\mathfrak{M}^{\prime})$. Set all constants to 0 and $R$ to be empty by default. We are left to describe the transition function of $\mathfrak{M}^{\prime}$, which we shall do by cases. In our description, $I$ denotes the input to $\mathfrak{M}^{\prime}$.
@@ 2. Case 2:
    $c_{0}\neq 0\wedge d\neq 0$”. We do the dovetailing. In other words, for each $\beta$ satisfying $\neg R(\beta)\wedge\beta\leq c_{0}$”, we use $\mathfrak{M}$ to simulate $M$ on input $\langle\{\beta\},I\rangle$ for up to $c_{0}$ steps, using $c_{1}$ and $c_{2}$ to keep track of the counts. At the end of every step in simulation, we check if $\mathfrak{M}$ terminates. If so, we set $R(\beta)$ to true and move on to simulate $M$ on input $\langle\{\beta+1\},I\rangle$, if $\beta\neq c_{0}$. At the end of all the simulations and possible updates to $R$, set $c_{1}$ and $c_{2}$ to 0 and increment $c_{0}$. The constant $d$ is left unchanged.
@@ 4. Case 4:
  2. (2)
    $c_{0}$ first returns to 0 in a state other than an initial state after exactly $\kappa+1$ many steps, at which point $M$ would have been simulated on every input in
@@
    for up to arbitrarily many steps $\kappa$, with all eventual haltings recorded in $R$.
@@
Lemmas 3.37 and 3.40 reflect the difference in computability between considering all terminating runs of GSeqAPs and considering only short terminating runs, for regular ordinal tape lengths. In tandem, the lemmas also give us a nice characterisation of the reducibility relation $\leq^{P}$.
Theorem 3.41.
@@
Apply Theorem 3.41 with $\emptyset$ in place of $A$. ∎
@@ 4 Discussion and Questions
  2. (GCT2)
    a “high-level” notion of abstract algorithms such as GSeqAPs, and
  3. (GCT3)
    an algebraic model of set generation such as relativised constructibility
@@
Just as practical implementations of finitary SeqAs do not aim to model algorithmic intuition beyond the limits of Church-Turing thesis, GSeqAPs should not allow for algorithms that are non-constructible. We can view GSeqAPs as a paradigm for easy high-level descriptions of algorithms on arbitrary sets, one that gives the user enough structure to organise their ideas, and serves as guardrails to rule out clearly non-algorithmic descriptions.
@@
[…]
@@
    $\mathsf{ZFC}+\text{``}\exists\text{ a strongly compact cardinal''}\vdash\text{``}V\neq L[A]\text{ for all }A\in V\text{''}$.
In particular, (2) of Fact 4.3 is implied by Exercise 20.2 of [12]. Note that $V\neq L[A]$ for all $A\in V$ is equivalent to $(\mathcal{D}_{P},\leq_{\mathcal{D}_{P}})$ having no maximal element, so Fact 4.3 actually gives us
@@
[…]
@@
[…]
@@
Let us now turn our attention to the bounded relations $\leq^{P,s}_{\alpha}$ and $\leq^{P}_{\alpha}$. Let $\leq_{\alpha}$, $\leq_{w\alpha}$ and $\preceq_{\alpha}$ respectively denote the relations $\alpha$-recursive in”, “weakly $\alpha$-recursive in” and $\alpha$-computable in”. The first two relations are well-studied in the once-bustling field of $\alpha$-recursion theory (see e.g. Part C of [17]), whereas the last relation comes from the emerging field of $\alpha$-computability theory. It is known (Section 3 of [14]) that
@@
for admissible ordinals $\alpha$. The last inclusion cannot be reversed in general because $\leq^{P}_{\alpha}$ is transitive for all admissible $\alpha$ but $\preceq_{\alpha}$ is not. On the other hand, it is clear that $\leq^{P,s}_{\alpha}\ \subset\ \leq^{P}_{\alpha}$ for all limit $\alpha$, and the reverse is not true due to Lemmas 3.37 and 3.40. Further, for admissible $\alpha$, $\leq_{w\alpha}\ \neq\ \leq^{P,s}_{\alpha}\ \neq\ \preceq_{\alpha}$ in general because $\leq^{P,s}_{\alpha}$ is transitive unlike the other two relations. We know $\leq^{P,s}_{\alpha}\ \neq\ \leq_{\alpha}$ in general because $\leq^{P,s}_{\alpha}$ is upward consistent in the sense of Proposition 3.12, but $\leq_{\alpha}$ is not by Theodore Slaman’s answer in [18].
Table 3 summarises the similarities and differences among the relations $\leq^{P,s}_{\alpha}$, $\leq^{P}_{\alpha}$, $\leq_{\alpha}$ and $\preceq_{\alpha}$. When we say a relation possesses the property of “oracle-analogue”, we mean that the definition of said relation depends on — or makes reference to — some object analogous to the oracle in the definition of classical oracle machines.
Table 3: Comparing two of our new relative computability relations against more standard relations in the study of generalised computability, at an arbitrary admissible ordinal $\alpha$.
@@
[…]